Calculate the number of formula units in 25.0 mol NaNO3.
1
Expert's answer
2020-10-22T08:31:43-0400

Solution.

"\frac{m}{M} = \frac{N}{Na}"

"N = 25.0 \times 6.02 \times 10^{23} = 1505 \times 10^{22}"

Answer:

"N = 1505 \times 10^{22}"